OpenStudy (anonymous):

Two lines, C and D, are represented by the equations given below: Line C: y = x + 14 Line D: y = 3x + 2 Which of the following shows the solution to the system of equations and explains why? (6, 20), because both lines pass through this point (6, 20), because the point does not lie on any axis (3, 11), because one of the lines passes through this point (3, 11), because the point lies between the two axes

ILovePuppiesLol (ilovepuppieslol):

vampire diraries?

OpenStudy (anonymous):

yes

OpenStudy (welshfella):

the solution is the coordinates of the points of intersection of the 2 lines (given that they do intersect, of course)

OpenStudy (anonymous):

please help!

OpenStudy (anonymous):

i dont understand help plz @welshfella

OpenStudy (isaidavila):

To find the intersection of two lines just match both equations x+14=3x+2 x=6 y=6+14=20

OpenStudy (isaidavila):

Because both lines lie in this point